# Human to Document, AI to Code: Three Case Studies of Comparing GenAI for Notebook Competitions

### v1

@@ 5.2. Threats to Validity
External Validity Three threats to external validity merit discussion.
@@
  • First, generalizability beyond Kaggle competitions may be limited. Our RQ1 findings that documentation features are the strongest discriminators between gold and non-gold medalists and RQ2 results showing human superiority in documentation accessibility versus GenAI superiority in code quality may not apply to industry contexts where performance metrics differ. However, these patterns likely extend beyond competitions as they reflect fundamental differences in how current GenAI systems generate code versus explanatory text.
@@
  • Second, temporal validity affects our GenAI methodology. The models used in RQ2 represent current capabilities but could be outdated in the future. We mitigate this by focusing on fundamental trade-offs between code standardization and human-readable explanation revealed across all three competitions.
@@
  • Third, It is acknowledged that the sample size of 9 GenAI-generated notebooks constitutes a limitation of this study. This number was determined by the exploratory nature of the research and the computational resources required for generation and analysis. While the findings provide initial insights, a larger and more diverse sample of GenAI notebooks is necessary to ensure the generalizability of the results. Future work should aim to expand this dataset significantly.
Construct Validity One threat to construct validity requires acknowledgment.
@@
  • Medal-worthy performance definition: Competition rankings may not capture all dimensions of exceptional data science work. We address this by using established Kaggle medal criteria, which represent community-validated standards for excellence in competitive data science, providing a reliable benchmark for our analysis.
6. Related Work
Prior research by Rule et al. (Rule et al., 2018) established that computational notebooks are key tools for exploration and explanation, with well-crafted narratives being crucial for collaboration. This is particularly relevant in collaborative ecosystems like Kaggle, where Hayashi et al. (Hayashi et al., 2021) found public notebooks are central to knowledge sharing. Reinforcing this, an analysis by Wang et al. (Wang et al., 2021) revealed that community-defined quality on Kaggle correlates more strongly with documentation and readability than with model performance alone. However, transitioning notebooks from exploration to production-ready artifacts remains a significant challenge, often due to difficulties in maintaining quality and structured workflows, as noted by Quaranta et al. (Quaranta, 2022). This focus on human-centric communication aligns with creativity research by Bangerl et al. (Bangerl et al., 2024), which highlights a powerful complementarity: humans excel at generating novel ideas, while GenAI is superior at providing detailed, structured elaboration.
While these studies establish the importance of human-centric documentation and the differing capacities of humans and AI, how this complementarity manifests within the specific domain of data science notebooks remains underexplored. This study, therefore, investigates this intersection by comparing the code and documentation practices of human and GenAI notebooks in a competitive setting.
7. Conclusion
@@
This study conducted a comparative analysis of human-written and GenAI notebooks within the context of competitive data science to explore the distinct strengths of each. The findings reveal a clear and complementary set of advantages. GenAI excels at producing code with high technical quality, featuring fewer code smells and less technical debt, while human notebooks, particularly medal-winning entries, are distinguished by their comprehensive and accessible documentation.
@@
The results suggest that while GenAI can achieve superior code cleanliness, top-tier performance in these competitions currently depends more on human-centric strengths like creative problem-solving, algorithmic innovation, and insightful documentation rather than on code quality alone. The implications of these findings are threefold: they may allow competition hosts to identify features of GenAI usage, guide competitors in how to best use GenAI assistants, and help researchers understand which skills are not easily amplified by current AI tools. Ultimately, this work provides a foundation for exploring the intersections of GenAI development and human creativity, pointing toward a future of hybrid approaches where GenAI handles routine code generation while humans drive the innovative problem-solving and narrative communication that define exceptional data science.

### v2

@@ VI-A Internal Validity
VI-B External Validity
Three threats to external validity merit discussion.
@@ VI-B External Validity
  • Generalizability beyond Kaggle competitions may be limited. Our RQ1 findings that documentation features are the strongest discriminators between gold and non-gold medalists and RQ2 results showing human superiority in documentation accessibility versus GenAI superiority in code quality may not apply to industry contexts where performance metrics differ. However, these patterns likely extend beyond competitions as they reflect fundamental differences in how current GenAI systems generate code versus explanatory text.
@@
  • Temporal validity affects our GenAI methodology. The models used in RQ2 represent current capabilities but could be outdated in the future. We mitigate this by focusing on fundamental trade-offs between code standardization and human-readable explanation revealed across all three competitions.
@@
  • We acknowledge that the sample size of 9 GenAI-generated notebooks constitutes a limitation of this study. This number was determined by the exploratory nature of the research and the computational resources required for generation and analysis. While the findings provide initial insights, a larger and more diverse sample of GenAI notebooks is necessary to ensure the generalizability of the results. Future work should aim to expand this dataset significantly.
VI-C Construct Validity
One threat to construct validity requires acknowledgment.
@@ VI-C Construct Validity
  • Medal-worthy performance definition: Our first construct, notebook quality, is proxied by Kaggle’s medal system (i.e., Bronze, Silver, Gold). This is a potential threat, as competition rankings may not capture all dimensions of exceptional data science work, such as long-term maintainability, real-world business impact, or deployment efficiency outside the competition environment. However, we mitigate this by using the established Kaggle medal criteria, which represent community-validated standards for excellence in competitive data science. This provides a reliable and holistic benchmark for notebook quality within the competitive data science context.
  • Practical Usability of Code Quality Metrics: Our initial analysis defined code quality using static metrics like code smells and technical debt, which was a valid concern regarding construct validity as it does not capture runnability. To address this and clearly bound our claims, we conducted a post-review execution validation. As reported in Section IV-B, this check confirmed that three of GenAI notebooks failed to run, despite their high static quality. This finding reinforces the limitation and shows that our ’code quality’ construct does not fully map to ’practical usability’.
VII Related Work
VII-A Studies on Computational Notebooks
Prior research has established computational notebooks as a principal paradigm for data scientists, valued for their support of two distinct roles: private exploration and public explanation [18]. As literate programming environments, they are designed to integrate executable code, narrative documentation, and visualizations within a single document [7, 3]. However, achieving narrative coherence is not an automatic outcome. Kery et al. [7] underscore the significant human effort involved, observing that data scientists find it a non-trivial task to clean up and curate their messy, exploratory code into coherent “stories” for presentation or sharing .
In collaborative settings, a well-constructed narrative is indispensable for communicating results and knowledge dissemination. This dynamic is particularly evident in competitive platforms like Kaggle. Wang et al. [22] demonstrated that community-defined quality (i.e., highly-voted notebooks) correlates more strongly with extensive documentation and high readability than with the notebook’s objective performance score on the leaderboard . Their analysis of these well-documented notebooks revealed nine distinct categories of documentation, including process descriptions, headlines for navigation, and explanations for analytical reasoning . This finding underscores the role of notebooks as vital communication artifacts, not merely as code containers.
Yet, this ideal of a clean, well-documented notebook is often at odds with the reality of data science workflows. The transition from exploration to explanation is fraught with pain points. Chattopadhyay et al. [3] cataloged numerous challenges that disrupt data scientists, including difficulties in setting up environments, managing dependencies, versioning exploratory code, and deploying notebooks to production. This difficulty contributes to the well-documented notebook-to-production gap [16].
A critical aspect of this gap is the widespread lack of reproducibility. A large-scale study of notebooks from biomedical publications by Samuel et al. found that the ”large majority” of notebooks could not be automatically executed. The primary cause was ”issues with the documentation of dependencies” , with most notebooks failing due to ModuleNotFoundError or ImportError . That same study noted a key correlation: notebooks with a low ratio of markdown-to-code cells were more likely to have exceptions, directly linking a lack of documentation to poor reproducibility . This challenge is precisely what Quaranta et al. [16] identified as the key blocker: deficiencies in code quality, structure, and maintainability prevent notebooks from being reliably integrated into production workflows.
VII-B Human and GenAI Complementarity
This emphasis on human-driven documentation, narrative, and high-level strategy aligns with findings on human-GenAI complementarity. Research indicates that humans demonstrate superior performance in areas requiring deep domain expertise, complex logical reasoning, and creative solutions. For example, Licorish et al. [10] found that humans performed better on tasks requiring in-depth domain knowledge, such as quantum optimization algorithms or debugging complex logic. GenAI, in contrast, specializes in the rapid generation of structured content and the management of repetitive, boilerplate code [1]. How this complementary relationship manifests specifically within computational notebooks remains underexplored.
The introduction of GenAI is prompting a re-conceptualization of the developer’s role. Rather than authoring code line-by-line, the human is repositioned as a “curator” [4] or as a “system orchestrator” responsible for providing high-level intent  [20, 11, 6]. In this paradigm, the human directs the overall strategy and validates the AI’s output, while the AI manages the low-level implementation. Consequently, human validation emerges as a critical function.
Our research builds upon this concept by quantitatively comparing how this role differentiation manifests in a competitive environment. We examine the balance between documentation (a human strength) and technical code quality (a GenAI strength). Our work is supported by findings from Molison et al. [12], which found that code generated by LLMs generally contains fewer bugs and requires less remediation effort. They also observed that fine-tuning, while effective at reducing high-severity blocker and critical bugs by shifting them to lower-severity categories, simultaneously degraded the model’s overall performance. Moreover, for complex, competition-level tasks, LLMs were found to introduce structural problems not present in human-authored code. Similarly, Licorish et al. [10] found that GPT-4 passed a higher percentage of functional test cases than human-written code. Cotroneo et al. [4] add to this, noting that AI code is generally simpler and more repetitive.
VIII Conclusion
@@
This study conducted a comparative analysis of human-written and GenAI-generated notebooks within the high-stakes context of competitive data science. Our goal was to explore the distinct, and often complementary, strengths and weaknesses inherent in both human and AI approaches. The findings reveal a clear dichotomy: GenAI excels at producing code with high static technical quality, consistently featuring significantly fewer code smells and less technical debt than human-written counterparts. This suggests a strong capability for adhering to programming standards and generating clean, standardized code.
However, this technical superiority comes with a critical caveat. Our execution checks revealed that high static quality does not translate to practical runnability, as several GenAI notebooks failed during execution due to fundamental data-handling and logic errors. In contrast, human-written notebooks, particularly the medal-winning entries, were distinguished by their comprehensive, accessible, and insightful documentation. This aligns with our RQ1 findings, confirming that narrative communication and clear explanations are a hallmark of high-quality human work in this domain.
The results strongly suggest that while GenAI can achieve superior code cleanliness, top-tier performance in these competitions currently depends more on human-centric strengths. These include creative problem-solving, domain-specific algorithmic innovation, and the ability to craft insightful documentation that communicates a clear analytical narrative—skills that code quality metrics alone fail to capture.
The implications of these findings are threefold and significant for the data science community:
  • For competition hosts: The identifiable features of GenAI-generated code (e.g., high static quality but low readability scores) may allow for the development of new heuristics to detect GenAI usage, which is crucial for upholding competition integrity and rules.
  • For competitors: These results provide a strategic guide for using GenAI assistants. Competitors can leverage these tools to handle routine tasks like code standardization and boilerplate generation, freeing up their own time to focus on high-impact, human-centric tasks such as feature engineering, model innovation, and narrative-building.
  • For researchers and educators: Our findings help identify which critical data science skills, such as narrative reasoning, complex problem decomposition, and practical validation, are not easily amplified or replicated by current AI tools, guiding future research and curriculum development.
@@
Ultimately, this work provides a foundation for exploring the intersections of GenAI development and human creativity. It points toward a future of hybrid approaches where GenAI handles routine code generation and quality assurance, while humans drive the innovative problem-solving and narrative communication that define exceptional data science.
IX Acknowledgement
This research is supported by JSPS Kakenhi (A) JP24H00692.
X Data Availability
To facilitate replication studies, our scripts and dataset are publicly available online, which can be found at